OpenStudy (cutiecomittee123):

Solve on the interval [0,2pi) 2sin^2x+3sinx+1=0

OpenStudy (cutiecomittee123):

@mathmate @satellite73 @freckles

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

set each factor equal to zero, solve for sine then solve for x

Nnesha (nnesha):

let sin(x) = y \[2y^2+3y+1=0\] can you factor the quadratic equation ?

OpenStudy (cutiecomittee123):

well i tried before but i can try again

Nnesha (nnesha):

nice try it let me know what youge :=))

Nnesha (nnesha):

t*

OpenStudy (cutiecomittee123):

x=-3+sqrt1/4 x=-3-sqrt1/4

OpenStudy (cutiecomittee123):

@Nnesha

Nnesha (nnesha):

do you mean \[x=\frac{ -3 \pm \sqrt{1} }{ 4 }\] ?

OpenStudy (cutiecomittee123):

Yeah i just wrote them out, same thing

Nnesha (nnesha):

i just want make sure it's (-3-sqrt{1}) over 4 or -3 - (sqrt{1}/4})

Nnesha (nnesha):

alright square rooot of 1 is just one \[\frac{ -3 +1 }{ 4 }~,~~\frac{-3-1}{4}\] simplify this brb let me login to other laptop

OpenStudy (cutiecomittee123):

okay

Nnesha (nnesha):

hmm there is a typo

OpenStudy (cutiecomittee123):

x=-1, -1/2

OpenStudy (cutiecomittee123):

right

Nnesha (nnesha):

yes that's right \[\frac{ -3+1 }{ 4 } \rightarrow \frac{-2}{4} \rightarrow \frac{-1}{2}\]

OpenStudy (cutiecomittee123):

sweet but how arethese answers relevant to the solution because i know that these arent the answers since my assignment has answers like 2pi/3

Nnesha (nnesha):

so sin(x) = -1/2 and -1 u can use unit circle to find exact solution both values solutions are negative sign chart  so answer would be in 3rd and 4th quadrant